OpenStudy (anonymous):

Solve (sqrt x - 2) = 5

OpenStudy (anonymous):

\[\sqrt x-2=5\]\[\sqrt x=7\]Now what do you think we should do to get x by itself

OpenStudy (anonymous):

Hint:\[(x^a)^b=x^{ab}\]\[\sqrt x =x^{{1 \over 2}}\]

OpenStudy (anonymous):

\[ \sqrt{x - 2} = 5\] this is how the equation is not what you just did

OpenStudy (anonymous):

\[\sqrt {x-2}=5\]Same principle applies. To get rid of a square root (^half power) we need to square both sides\[x-2=25\]

OpenStudy (anonymous):

\[\sqrt {x-2}=(x-2)^{{1 \over 2}}\]\[((x-2)^{{1 \over 2}})^2=(x-2)^{({1 \over 2}\times 2)}=x-2\]

OpenStudy (anonymous):

so would this be a extraneous solution?

OpenStudy (anonymous):

I don't know what that means but there is a solution

OpenStudy (anonymous):

okay its when you put the solution into you equation and see if it equals , so in this situation 5 so i should put -2 into x to see if it will equal to 5 right?

OpenStudy (anonymous):

\[\sqrt {x-2}=5\]\[(\sqrt {x-2})^2=5^2\]\[x-2=25\]\[x=27\]

OpenStudy (anonymous):

I was never good at remembering def'n only the math. idk

OpenStudy (anonymous):

okay well thanks tho